OpenStudy (anonymous):

For the limit lim x → 2 (x3 − 4x + 9) = 9 illustrate the definition by finding the largest possible values of δ that correspond to ε = 0.2 and ε = 0.1

OpenStudy (anonymous):

We can confirm the limit:\[ \lim_{x\to a} f(x) = L \]If we can show that for any \(\epsilon\) there exists a \(\delta\) such that:\[ \forall x:\quad|x-a| < \delta \implies |f(x) - L| < \epsilon \]

OpenStudy (anonymous):

after i take the 9 to right side and subtract it i get 0 i take x common then what ?

OpenStudy (anonymous):

Let's start with \(\epsilon =0.2\)

OpenStudy (anonymous):

alright then ?

OpenStudy (anonymous):

\[ |(x^2-4x+9)-9|<0.2 \]

OpenStudy (anonymous):

You need to solve for \(x\) here.

OpenStudy (anonymous):

alrights lets see

OpenStudy (anonymous):

\[ |x^2-4x|<0.2 \]

OpenStudy (anonymous):

\[ |a\cdot b|=|a||b| \]

OpenStudy (anonymous):

\[ |x||x-4|<0.2 \]

OpenStudy (anonymous):

i make it \[x ^{3}-4x\]

OpenStudy (anonymous):

Oh right... sorry.

OpenStudy (anonymous):

\[ |x^3-4|<0.2\implies -0.2<x^3-4x<0.2 \]

OpenStudy (anonymous):

hmm that seems rigth

OpenStudy (anonymous):

For the \(\delta\) we find:\[ -\delta <x-2<\delta \]

OpenStudy (anonymous):

By the way \(x^3-4x=x(x^2-4)=x(x-2)(x+2)\)

OpenStudy (anonymous):

yes i was thinkiong of that . but couldnt solve after it

OpenStudy (anonymous):

What we basically have is: \[ \delta = \frac {0.2}{x(x+2)} \]

OpenStudy (anonymous):

If we let \(x=2\) at this point, then: \[ \delta = \frac{0.2}{2(2+2)}=\frac{0.2}{8} \]

OpenStudy (anonymous):

so now to slove for delta

OpenStudy (anonymous):

y we taking x as 2?

OpenStudy (anonymous):

Its \[ \lim_{x\to 2} \]

OpenStudy (anonymous):

my bad !

OpenStudy (anonymous):

Though we should definitely double check our answer now.

OpenStudy (anonymous):

i have lost all hopes with this question

OpenStudy (anonymous):

alright let me ask someone

OpenStudy (anonymous):

Okay sorry, we did something wrong.

OpenStudy (anonymous):

no one to confirm bro . but if we solving for 0.1 we subsitute 0.2 right?

OpenStudy (anonymous):

did we?

OpenStudy (anonymous):

Yeah, I shouldn't have done \(x=2\).

OpenStudy (anonymous):

but its tending toward 2 right?

OpenStudy (anonymous):

Yeah, but at \(x=2\) we underestimate the epsilon.

OpenStudy (anonymous):

Okay, I got another idea!

OpenStudy (anonymous):

hmmm . my friend got the first one as 0.01

OpenStudy (anonymous):

That is not the largest possible \(\delta\) though.

OpenStudy (anonymous):

hmmm shes gonna tell the method then maybe it could help us

OpenStudy (anonymous):

My next solution is to try letting \(x=2+\delta\)

OpenStudy (anonymous):

\[ (2+\delta)^3-4(2+\delta)-9=9+0.2 \]

OpenStudy (anonymous):

\[6\delta ^{2} + 8\delta+\delta ^{3} - 9\]

OpenStudy (anonymous):

= 9+0.2

OpenStudy (anonymous):

that is correct if we solve it

OpenStudy (anonymous):

now lets see what my friend says

OpenStudy (anonymous):

\[ \delta<1.13219\\ \delta<1.11637 \]

OpenStudy (anonymous):

so its 1.1

OpenStudy (anonymous):

They both must be true, so that means \(1.11637\) is the largest.

OpenStudy (anonymous):

hmmm i see and for 0.1 we just add that pellet?

OpenStudy (anonymous):

to the right side*

OpenStudy (anonymous):

Oh no, I typed it in wrong before!

OpenStudy (anonymous):

damn :P

OpenStudy (anonymous):

now we got 3 solutions

OpenStudy (anonymous):

Yeah, you always take the smallest one.

OpenStudy (anonymous):

I mean the one that is closes to 0.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%282%2B0.024%29%5E3%E2%88%924%282%2B0.024%29%2B9%3D9%2Bepsilon Look at how close \(|\delta| =0.024\) is!

OpenStudy (anonymous):

Yeah, I'd go with \[ \delta \approx \pm 0.025485 \]

OpenStudy (anonymous):

sweet

OpenStudy (anonymous):

It's the lowest of all 6 deltas.

OpenStudy (anonymous):

hmm true

OpenStudy (anonymous):

u helped a lottt broo

OpenStudy (anonymous):

When epsilon is 0.1 the lowest delta we have is: 0.0123847

OpenStudy (anonymous):

yep saw that

OpenStudy (anonymous):

Those are pretty accurate.

OpenStudy (anonymous):

I feel bad using wolfram for this, but cubic roots is NOT a trivial thing.

OpenStudy (anonymous):

The cubic formula looks like a monster that ate ten quadratic formulas.

OpenStudy (anonymous):

hahhaha thats one way to put it:P

OpenStudy (anonymous):

See for yourself.

OpenStudy (anonymous):

its not hard bt this question is the first one i encountered trouble with

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."